OpenStudy (anonymous):

4k+2(2^3)+2k greater than equal to 52+k-6

OpenStudy (anonymous):

\[4k+2^{3}+2k \ge 52+k+6\] \[5k+8 \ge 58\] -8 -8 5k=50 k=10